OpenStudy (anonymous):

What are the solutions of 3x^2 + 6x + 6 = 0? Answer A. 1 ± i B. –1 ± i C. –2 ± 2i D. –2 ± i What are the solutions of 3x^2 + 6x + 6 = 0? Answer A. 1 ± i B. –1 ± i C. –2 ± 2i D. –2 ± i @Mathematics

OpenStudy (anonymous):

SIMPLIFY: x^2 + 2x + 2 = 0 x = -2 +-sqrt(4 - 4*1*2) / 2

OpenStudy (anonymous):

divide by 3 first to start with \[x^2+2x+2=0\] then \[x^2+2x=-2\] \[(x+1)^2=-2+1=-1\] \[x+1=\pm\sqrt{-1}\] \[x+1=\pm i\] \[x=-1\pm i\]

OpenStudy (anonymous):

Thank you!!